# Fortune: Formula-Driven Reinforcement Learning for Symbolic Table Reasoning in Language Models

### v2

@@ Appendix C Supplementary Proofs
C.1 Proof of Symbolic Reasoning Dominance
@@
By Assumption 1 (3), both the symbolic policy $\pi_{\theta}^{\mathrm{sym}}$ and the textual policy $\pi_{\theta}^{\mathrm{txt}}$ are equally capable of producing such high-level plans:
@@
We now compare how these two policies operationalize the same plan downstream:
  • Symbolic reasoning. After selecting a correct high-level plan, the symbolic policy proceeds by emitting a formal expression—typically a spreadsheet formula $f$—that directly encodes the solution. This formula is then passed to an external executor, which deterministically computes the final answer $a=\operatorname{exec}(f,\mathbb{T})$. Under Assumptions 1 (1) and (2), if the plan is correct, the execution will reliably yield the correct answer $a^{\star}(s)$. Thus, the expected reward under the symbolic policy is:
@@
C.2 Proof of MLE Minimizes KL Divergence
@@
C.3 Proof of Convergence of SFT and Reward Upper Bound
@@
$\mathbb{E}_{s\sim p,f\sim\pi_{\theta^{\star}}}\left[r(\operatorname{exec}(f,\mathbb{T})\mid s)\right]=\mathbb{E}_{s\sim p,f\sim\pi_{g}}\left[r(\operatorname{exec}(f,\mathbb{T})\mid s)\right].$ (25)
@@
C.4 Proof of RL Superiority
Proof.
$\mathbb{E}_{s\sim p,\,f\sim\pi_{\theta^{\star}}}\left[r(\operatorname{exec}(f,\mathbb{T})\mid s)\right]=\mathbb{E}_{s\sim p,\,f\sim\pi_{g}}\left[r(\operatorname{exec}(f,\mathbb{T})\mid s)\right].$ (26)
$\mathbb{E}_{s\sim p,\,f\sim\pi_{\theta}^{\mathrm{RL}}}\left[r(\operatorname{exec}(f,\mathbb{T})\mid s)\right]>\mathbb{E}_{s\sim p,\,f\sim\pi_{g}}\left[r(\operatorname{exec}(f,\mathbb{T})\mid s)\right],$ (27)
which implies the desired result. ∎
@@ Appendix D Supplementary Discussion of Methodology
In addition to the formal analysis in Section 3.2, we highlight several conceptual advantages of symbolic reasoning for table understanding:
@@
We also expand upon the discussion in Section 3.3, comparing supervised fine-tuning (SFT) and reinforcement learning (RL) for symbolic reasoning:
@@ D.3 Practical Challenges of Formula Tuning
While reinforcement learning (RL) offers significant advantages for symbolic table reasoning, it also introduces several practical challenges, especially under the assumptions outlined in Section 3.2 and 3.3.
@@ Appendix E Detailed Settings of Experiments
Models. Our experiments include both open-source and proprietary models. For open-source models, we use Qwen2.5-Coder7B (Qwen2.5-Coder-7B-Instruct, Apache 2.0 License) [75] and LLaMA-3.18B (LLaMA-3.1-8B-Instruct, Meta Llama 3 Community Licence) [76]. For proprietary models, we evaluate OpenAI’s [^2] https://openai.com/policies/row-terms-of-use/ GPT-4o (gpt-4o-2024-11-20), GPT-4o-mini (gpt-4o-mini-2024-07-18), and O1 (o1-2024-12-17) as baselines.
Datasets. As shown in Table 3, we conduct experiments on seven diverse table understanding benchmarks: WikiTQ [61], TabFact [62], FinQA [63], HiTab [25], MultiHiertt [64], AIT-QA [65], and TableBench [66]. These datasets vary in domain coverage, table structures, and question complexity, collectively spanning the full spectrum of table understanding tasks. For MultiHiertt, which contains multiple tables, we concatenate them vertically to form a single spreadsheet-like table. For training, we combine the first five datasets into a unified training corpus and train the model jointly on this merged set. Each dataset is then evaluated individually. All original training and test splits are preserved, except for TabFact, from which we randomly sample 10,000 examples to prevent its abundance of relatively simple binary QA examples from dominating or skewing the training. Among these benchmarks, AIT-QA and TableBench are considered out-of-distribution (OOD) evaluation sets, while the remaining datasets are treated as in-distribution (ID). The characteristics of each dataset are summarized below:
@@
In-Distribution WikiTQ [61] 13,753 4,217 Relational Wikipedia CC-BY-SA-4.0 Link
TabFact [62] 10,000 2,024 Relational Wikipedia CC-BY-4.0 Link
FinQA [63] 6,251 1,147 Relational Finance MIT Link
HiTab [25] 7,399 1,583 Hierarchical Statistical Reports C-UDA 1.0 Link
MultiHiertt [64] 7,795 1,038 Multiple & Hierarchical Finance MIT Link
Out-of-Distribution AIT-QA [65] 515 Hierarchical Airline CDLA-Sharing-1.0 Link
TableBench [66] 883 Relational Cross Domain CC0-1.0 Link
Table Encoding. We adopt a table encoding method similar to SpreadsheetEncoder [21], which converts a table into a linearized markdown-style format. Each cell is represented by its spreadsheet address and value, forming text sequences such as A1,Year|A2,Profit. This encoding preserves both structural and content information, enabling the model to better understand cell-level references.
@@
To encourage adherence to this format, we introduce a lightweight format reward. If the output fails to follow the required structure (e.g., malformed tags or unparseable JSON), the model receives a penalty of $-2$. If the format is valid and the answer can be successfully parsed from the JSON object, a small positive reward of $+0.1$ is added to the answer-level reward. Therefore, the final reward is as:
@@
Baselines. We compare our proposed framework against a broad range of strong baselines, including both prompting-based and fine-tuning-based methods. To ensure a fair comparison, we require all methods to output short, deterministic answers rather than open-ended free-form text. Following this criterion, we exclude TableLLM [73], which relies on a critique model for answer evaluation and does not produce a directly verifiable answer string. Prompting-based methods currently dominate the TableQA landscape, with most relying on large closed-source models for performance. We compare Fortune and Fortune++ with several representative methods in this category: Binder [36], Dater [34], API-Assisted [11], Chain-of-Table [35], ReAcTable [67], Norm-DP [68], TIDE [69], E5 [70], and SS-CoT [71]. We also include TableMaster [11], a recent recipe-based prompting framework built on GPT-4o-mini. For fine-tuning-based methods, we select models specifically trained for table question answering tasks. These include TAPEX-Large [31], OmniTab [72], TableLlama [32], TableGPT2 [74], and TabAF [22], a recent strong method that combines formula generation and hybrid self-consistency. Our framework is evaluated under the same settings to ensure consistency and comparability across methods.
@@
RL Training. We use Proximal Policy Optimization (PPO) for reinforcement learning (RL). The maximum prompt length is set to 8192 tokens, and the maximum response length is 512 tokens. The critic model is initialized with the same weights as the actor model. The actor is trained with a learning rate of 1e-6, while the critic uses a slightly higher learning rate of 1e-5 to enable faster value estimation. We set the KL divergence coefficient to 0.001 to balance exploration and policy stability. The generation temperature is set to 0.6 to encourage a mix of determinism and diversity in the generated reasoning chains and formula outputs. The PPO mini-batch size is 64. We evaluate performance every 20 steps and report the results based on the best performance achieved on each dataset.
@@
SFT Training. The supervised fine-tuning (SFT) training corpus is distilled from GPT-4o by prompting it with ground-truth answers, eliciting chain-of-thought reasoning followed by a final answer. We adopt a rejection-based fine-tuning (RFT) strategy [77], retaining only examples where the generated answer exactly matches the ground truth. For symbolic reasoning tasks, correctness is determined by executing the generated formula and verifying that the resulting answer matches the expected output. This approach ensures high-quality supervision for fine-tuning. All SFT models are trained for 4 epochs, and we report results based on the checkpoint with the highest exact match accuracy on the test set. We use a learning rate of 2e-5 and a batch size of 64.
@@
Evaluation Inference. For all models, including both open-source and proprietary ones, we use a temperature of 0, top-$k$ of 50, and top-$p$ of 0.7 during inference. Setting the temperature to 0 encourages deterministic outputs and improves stability in single-pass predictions. For self-consistency decoding in Fortune++, we use a higher temperature of 0.6 to promote diversity across multiple samples, enabling the model to better explore reasoning variations and improve final answer voting.
@@
Evaluation Metrics. Following prior work [61, 25], we primarily use exact match (EM) as the evaluation metric, applying numeric tolerance when comparing numerical values. Official evaluation scripts are used whenever available to ensure consistency. For TabFact, which is formulated as a binary classification task, we report standard classification accuracy. Since our training objective aligns with evaluation, we also use exact match (EM) for answer reward calculation.
@@
Software. We implement Fortune using Python 3.11, with the VERL framework [78] serving as the core architecture for reinforcement learning and other supervised fine-tuning with language models. Our implementation utilizes VLLM (v0.8.3) [79] for efficient LLM inference and generation, PyTorch (v2.4.0) with CUDA 12.8 for deep learning operations, and Ray [80] for distributed training and inference. FlashAttention-2 [81] is integrated to accelerate attention computation. For proprietary LLMs, we access OpenAI models via the Microsoft Azure platform [^3] https://azure.microsoft.com/. For formula execution, we use the open-source spreadsheet engine formulas [^4] https://github.com/vinci1it2000/formulas (EUPL 1.1+ License), which supports a wide range of standard spreadsheet operators. A representative list of symbolic operators used in our table reasoning framework is provided in Appendix K.
@@
Appendix F Ablation Study and Upper-Bound Performance of Fortune++
Table 4: Performance comparison of TabAF, Fortune, and Fortune++ variants. Values in the table indicate accuracy (%). ‘-’ indicates results not reported in the related paper. Gray rows represent upper-bound performance. Numbers in parentheses with a downward arrow ($\downarrow$) indicate the performance drop relative to the default Fortune++ configuration. Top results are highlighted in dark blue.

### v3

@@ Appendix C Supplementary Proofs
C.1 Proof of Symbolic Reasoning Superiority
Lemma 1 (Reward Decomposition).
For textual reasoning, the expected reward is:
$\scriptstyle\mathbb{E}_{f\sim\pi_{\theta}^{\mathrm{sym}}}[r(\operatorname{exec}(f,\mathbb{T})\mid s)]=\sum_{f}\pi_{\theta}^{\mathrm{sym}}(f\mid s)\cdot\mathbbm{1}[\operatorname{exec}(f,\mathbb{T})=a^{\star}(s)].$ (9)
Assumption 1 (Symbolic Reasoning Setting).
  1. 1.
  2. 2.
  3. 3.
    Both textual and symbolic policies are assumed capable of representing valid high-level solution strategies in their respective formats, namely text or formula.
@@ C.1 Proof of Symbolic Reasoning Superiority
By Assumption 1 (3), both the symbolic policy $\pi_{\theta}^{\mathrm{sym}}$ and the textual policy $\pi_{\theta}^{\mathrm{txt}}$ are assumed capable of producing such high-level plans:
@@
We now compare how these two policies execute the same plan downstream.
  • Symbolic reasoning. After selecting a correct high-level plan, the symbolic policy proceeds by emitting a formal expression—typically a spreadsheet formula ff—that directly encodes the solution. This formula is then passed to an external executor, which deterministically computes the final answer $a=\operatorname{exec}(f,\mathbb{T})$. Under Assumptions 1 (1) and (2), if the plan is correct, the execution will reliably yield the correct answer $a^{\star}(s)$. Thus, the expected reward under the symbolic policy is:
@@
C.2 Proof of RL Superiority
Assumption 2 (SFT Setting).
  1. 1.
    Sufficient expressivity. The model class $\{\pi_{\theta}(\cdot\mid s)\}$ is expressive enough to represent the teacher policy $\pi_{g}(\cdot\mid s)$ (a stronger model, e.g., GPT-4o), in the sense that
  2. 2.
  3. 3.
Lemma 2 (MLE Minimizes KL Divergence).
$\scriptstyle\mathbb{E}_{a\sim\pi_{g}}\left[-\log\pi_{\theta}(a\mid s)\right]=H\left(\pi_{g}(\cdot\mid s)\right)+D_{\mathrm{KL}}\left(\pi_{g}\parallel\pi_{\theta}\right),$ (15)
The proof of Lemma 2 is provided in Appendix C.3.
Lemma 3 (Convergence of SFT and Reward Upper Bound).
Under Assumption 2, the optimal supervised fine-tuning (SFT) policy
satisfies
$\scriptstyle\mathbb{E}_{s\sim p,\,f\sim\pi_{\theta^{\star}}}\left[r(\operatorname{exec}(f,\mathbb{T})\mid s)\right]\leq\mathbb{E}_{s\sim p,\,f\sim\pi_{g}}\left[r(\operatorname{exec}(f,\mathbb{T})\mid s)\right].$ (18)
The proof of Lemma 3 is provided in Appendix C.4.
Remark 3 (SFT Bound).
Assumption 3 (RL Exploration).
For each input $s\in\mathcal{S}$, we assume that the policy distribution $\pi_{\theta}(\cdot\mid s)$ assigns non-zero probability mass to at least one correct action with reward $r(a^{\star}\mid s)=1$. This does not require the policy to sample a correct action at every step, only that the support of the distribution includes some high-reward actions, so they may be discovered over the course of training.
Proof.
$\scriptstyle\mathbb{E}_{s\sim p,\,f\sim\pi_{\theta^{\star}}}\left[r(\operatorname{exec}(f,\mathbb{T})\mid s)\right]=\mathbb{E}_{s\sim p,\,f\sim\pi_{g}}\left[r(\operatorname{exec}(f,\mathbb{T})\mid s)\right].$ (19)
$\scriptstyle\mathbb{E}_{s\sim p,\,f\sim\pi_{\theta}^{\mathrm{RL}}}\left[r(\operatorname{exec}(f,\mathbb{T})\mid s)\right]>\mathbb{E}_{s\sim p,\,f\sim\pi_{g}}\left[r(\operatorname{exec}(f,\mathbb{T})\mid s)\right],$ (20)
which implies the desired result. ∎
C.3 Proof of MLE Minimizes KL Divergence
@@
C.4 Proof of Convergence of SFT and Reward Upper Bound
@@
$\scriptstyle\mathbb{E}_{s\sim p,f\sim\pi_{\theta^{\star}}}\left[r(\operatorname{exec}(f,\mathbb{T})\mid s)\right]=\mathbb{E}_{s\sim p,f\sim\pi_{g}}\left[r(\operatorname{exec}(f,\mathbb{T})\mid s)\right].$ (27)
@@ Appendix D Supplementary Discussion of Methodology
In addition to the formal analysis in Section B.1, we highlight several conceptual advantages of symbolic reasoning for table understanding:
@@
We also expand upon the discussion in Section B.2, comparing supervised fine-tuning (SFT) and reinforcement learning (RL) for symbolic reasoning:
@@ D.3 Practical Challenges of Formula Tuning
While reinforcement learning (RL) offers significant advantages for symbolic table reasoning, it also introduces several practical challenges, especially under the assumptions outlined in Section B.1 and B.2.
@@ Appendix E Detailed Settings of Experiments
Models. Our experiments include both open-source and proprietary models. For open-source models, we use Qwen2.5-Coder7B (Qwen2.5-Coder-7B-Instruct, Apache 2.0 License) (Hui et al., 2024) and LLaMA-3.18B (LLaMA-3.1-8B-Instruct, Meta Llama 3 Community Licence) (AI, 2024).
Datasets. As shown in Table 4, we conduct experiments on seven diverse table understanding benchmarks: WikiTQ (Pasupat and Liang, 2015), TabFact (Chen et al., 2020), FinQA (Chen et al., 2021b), HiTab (Cheng et al., 2021), MultiHiertt (Zhao et al., 2022), AIT-QA (Katsis et al., 2021), and TableBench (Wu et al., 2025). These datasets vary in domain coverage, table structures, and question complexity, collectively spanning the full spectrum of table understanding tasks. For MultiHiertt, which contains multiple tables, we concatenate them vertically to form a single spreadsheet-like table. For training, we combine the first five datasets into a unified training corpus and train the model jointly on this merged set. Each dataset is then evaluated individually. All original training and test splits are preserved, except for TabFact, from which we randomly sample 10,000 examples to prevent its abundance of relatively simple binary QA examples from dominating or skewing the training. Among these benchmarks, AIT-QA and TableBench are considered out-of-distribution (OOD) evaluation sets, while the remaining datasets are treated as in-distribution (ID). The characteristics of each dataset are summarized below:
@@
In-Distribution WikiTQ (Pasupat and Liang, 2015) 13,753 4,217 Relational Wikipedia $C$-BY-SA-4.0 Link
TabFact (Chen et al., 2020) 10,000 2,024 Relational Wikipedia $C$-BY-4.0 Link
FinQA (Chen et al., 2021b) 6,251 1,147 Relational Finance MIT Link
HiTab (Cheng et al., 2021) 7,399 1,583 Hierarchical Statistical Reports C-UDA 1.0 Link
MultiHiertt (Zhao et al., 2022) 7,795 1,038 Multiple & Hierarchical Finance MIT Link
Out-of-Distribution AIT-QA (Katsis et al., 2021) 515 Hierarchical Airline CDLA-Sharing-1.0 Link
TableBench (Wu et al., 2025) 883 Relational Cross Domain CC0-1.0 Link
Table Encoding. We adopt a table encoding method similar to SpreadsheetEncoder (Dong et al., 2024), which converts a table into a linearized markdown-style format. Each cell is represented by its spreadsheet address and value, forming text sequences such as A1,Year|A2,Profit. This encoding preserves both structural and content information, enabling the model to better understand cell-level references.
@@
To encourage adherence to this format, we introduce a lightweight format reward. If the output fails to follow the required structure (e.g., malformed tags or unparseable JSON), the model receives a penalty of 2-2. If the format is valid and the answer can be successfully parsed from the JSON object, a small positive reward of +0.1+0.1 is added to the answer-level reward. Therefore, the final reward is as:
@@
Baselines. We compare our proposed framework against a broad range of strong baselines, including both prompting-based and fine-tuning-based methods. To ensure a fair comparison, we require all methods to output short, deterministic answers rather than open-ended free-form text. Following this criterion, we exclude TableLLM (Zhang et al., 2025), which relies on a critique model for answer evaluation and does not produce a directly verifiable answer string. Prompting-based methods currently dominate the TableQA landscape, with most relying on large closed-source models for performance. We compare Formula-R1 and Formula-R1-Plus with several representative methods in this category: Binder (Cheng et al., 2023), Dater (Ye et al., 2023), API-Assisted (Cao and Liu, 2025), Chain-of-Table (Wang et al., 2024), ReAcTable (Zhang et al., 2023), Norm-DP (Liu et al., 2023), TIDE (Yang et al., 2025), E5 (Zhang et al., 2024c), and SS-CoT (Zhao et al., 2024b). We also include TableMaster (Cao and Liu, 2025), a recent recipe-based prompting framework built on GPT-4o-mini. For fine-tuning-based methods, we select models specifically trained for table question answering tasks. These include TAPEX-Large (Liu et al., 2022), OmniTab (Jiang et al., 2022), TableLlama (Zhang et al., 2024a), TableGPT2 (Su et al., 2024), and TabAF (Wang et al., 2025), a recent strong method that combines formula generation and hybrid self-consistency. Our framework is evaluated under the same settings to ensure consistency and comparability across methods.
@@
RL Training. We use Proximal Policy Optimization (PPO) for reinforcement learning (RL). The maximum prompt length is set to 8192 tokens, and the maximum response length is 512 tokens. The critic model is initialized with the same weights as the actor model. The actor is trained with a learning rate of 1e-6, while the critic uses a slightly higher learning rate of 1e-5 to enable faster value estimation. We set the KL divergence coefficient to 0.001 to balance exploration and policy stability. The generation temperature is set to 0.6 to encourage a mix of determinism and diversity in the generated reasoning chains and formula outputs. The PPO mini-batch size is 64.
@@
SFT Training. The supervised fine-tuning (SFT) training corpus is distilled from GPT-4o by prompting it with ground-truth answers, eliciting chain-of-thought reasoning followed by a final answer. We adopt a rejection-based fine-tuning (RFT) strategy (Yuan et al., 2023), retaining only examples where the generated answer exactly matches the ground truth. For symbolic reasoning tasks, correctness is determined by executing the generated formula and verifying that the resulting answer matches the expected output. This approach ensures high-quality supervision for fine-tuning. All SFT models are trained for 4 epochs, and we report results based on the checkpoint with the highest exact match accuracy on the test set. We use a learning rate of 2e-5 and a batch size of 64.
@@
Evaluation Inference. For all models, including both open-source and proprietary ones, we use a temperature of 0, top-$k$ of 50, and top-$p$ of 0.7 during inference. Setting the temperature to 0 encourages deterministic outputs and improves stability in single-pass predictions. For self-consistency decoding in Formula-R1, we use a higher temperature of 0.6 to promote diversity across multiple samples, enabling the model to better explore reasoning variations and improve final answer voting.
@@
Evaluation Metrics. Following prior work (Pasupat and Liang, 2015; Cheng et al., 2021), we primarily use exact match (EM) as the evaluation metric, applying numeric tolerance when comparing numerical values. Official evaluation scripts are used whenever available to ensure consistency. For TabFact, which is formulated as a binary classification task, we report standard classification accuracy. Since our training objective aligns with evaluation, we also use exact match (EM) for answer reward calculation.
@@
Software. We implement Fortune using Python 3.11, with the VERL framework (Sheng et al., 2024) serving as the core architecture for reinforcement learning and other supervised fine-tuning with language models. Our implementation utilizes VLLM (v0.8.3) (Kwon et al., 2023) for efficient LLM inference and generation, PyTorch (v2.4.0) with CUDA 12.8 for deep learning operations, and Ray (Moritz et al., 2018) for distributed training and inference. FlashAttention-2 (Dao et al., 2022) is integrated to accelerate attention computation. For proprietary LLMs, we access OpenAI models via the Microsoft Azure platform [^1] https://azure.microsoft.com/. For formula execution, we use the open-source spreadsheet engine formulas [^2] https://github.com/vinci1it2000/formulas (EUPL 1.1+ License), which supports a wide range of standard spreadsheet operators. A representative list of symbolic operators used in our table reasoning framework is provided in Appendix K.
@@
Appendix F Reward Sensitivity Analysis
In this section, we further analyze the sensitivity of Formula Tuning to the reward design. In particular, we focus on the partial reward assigned to executable but incorrect formulas. This intermediate reward is introduced to alleviate reward sparsity. In preliminary experiments, using only binary rewards $\{1,0\}$ resulted in unstable optimization and often caused training to collapse to trivial outputs. A key reason is that small pretrained language models typically lack sufficient prior knowledge of spreadsheet-formula syntax, making it difficult to reliably optimize answer correctness before the model first learns to produce valid and executable formulas.
[…]
Partial reward (exec-but-wrong) 0.0 0.1 0.2 0.3 0.4
Accuracy 57.98 70.64 72.32 72.04 70.53
Several observations can be drawn from these results. First, the model is highly sensitive to removing partial credit entirely: when the reward for executable-but-incorrect outputs is set to 0.0, performance drops substantially to 57.98. This confirms that a purely binary reward is insufficient for stable optimization in symbolic table reasoning. Second, performance remains strong and relatively stable when the partial reward lies within a moderate range of 0.1 to 0.3, with the best result achieved at 0.2. This indicates that the method is not overly sensitive to the exact value of the partial reward, as long as a reasonable amount of intermediate credit is provided. Rather than biasing the model toward incorrect outputs, the partial reward serves as a meaningful intermediate learning signal that encourages exploration of valid formulas. Third, assigning an excessively large partial reward (e.g., 0.4) leads to a slight performance drop, suggesting that overly rewarding executability without sufficient emphasis on correctness may weaken the model’s incentive to optimize toward the final answer.
Overall, these findings support the reward design in Formula Tuning. The intermediate reward for executable-but-incorrect formulas improves optimization stability and helps the model acquire symbolic structure before answer correctness is fully optimized. At the same time, the results suggest that the method is fairly robust to the precise choice of this reward within a reasonable range. This design is particularly important in our setting, where valid formula generation itself is a nontrivial capability for small language models.
Appendix G Ablation Study and Upper-Bound Performance of Formula-R1
Table 6: Performance comparison of TabAF, Formula-R1, and Formula-R1-Plus variants. Values in the table indicate accuracy (%). ‘-’ indicates results not reported in the related paper. Gray rows represent upper-bound performance. Numbers in parentheses with a downward arrow ($\downarrow$) indicate the performance drop relative to the default Formula-R1-Plus configuration. Top results are highlighted in dark blue.